A taho vendor has to carry his wares on a 1.50 – m uniform pole. A 10-kg load

is suspended on one end and on the other end is a 15-kg load. How far from

the 15-kg load shall he carry his wares?


1
Expert's answer
2021-03-31T09:38:30-0400
"m_1gx=m_2g(l-x)\\15x=10(1.5-x)\\x=0.6\ m"